# The Unified Cognitive Consciousness Theory for Language Models: Semantic Anchoring, Threshold Activation, and Emergent Intelligence

### v3

@@ Introduction
With no semantic anchors, all four LLMs activate the dominant arithmetic pattern with highest density $\rho_{d}(P_{\text{arithmetic}})$.
Two-Shot Pattern Redefinition:
Example 1: 2 - 3 = 5  Example 2: 7 - 4 = 11
Question: 15 - 8 = ?  Answer: 23  (100% reinterpretation as addition)
Despite millions of associations of “--” with subtraction, just two examples achieve complete pattern redefinition. This validates Equation 1: the anchoring context $\mathcal{A}$ provides sufficient evidence to shift the posterior $p(P|\mathcal{A},C)$ from subtraction to addition patterns. Success occurs because the representation gap $d_{r}(P_{\text{add}},T)$ is minimal—both examples unambiguously demonstrate addition.
Threshold Boundary Detection:
When examples are ambiguous (33 - 27 = 60, 11 - 9 = 20), models diverge near the critical threshold $\alpha_{c}$:
  • Claude: Interprets as $|a-b|\times 10$ (accuracy: 240)
  • GPT-4o: Interprets as $(a-b)\times 10$ (accuracy: -240)
  • DeepSeek/Gemini: Interpret as $a+b$ (accuracy: 138)
This demonstrates threshold-crossing behavior: when $S(\mathcal{A})\approx\alpha_{c}$, small differences in model-specific pattern densities produce qualitatively different interpretations. (detailed analysis in Appendix B).
Case 1.2: Visual Pattern Recognition
Figure 1 demonstrates cross-modal validation of threshold-crossing dynamics. The 4‐shot anchoring mechanism follows $\mathsf{UCCT}$’s mathematical framework: pre‐trained pattern repository provides unconscious substrate, four labeled cat images determine pattern density $\rho_{d}(P_{\textit{cat}})$, and novel query produces semantic distance $d_{r}(P_{\text{cat}},T)$. The anchoring strength
$S(\mathcal{A})=\alpha\,\rho_{d}(P_{\textit{cat}})-\beta\,d_{r}(P_{\text{cat}},T)-\gamma\log 4$
exceeds the critical threshold $\alpha_{c}$, triggering phase transition from random pattern activation to reliable cat classification. This validates $\mathsf{UCCT}$’s universality: identical mathematical principles govern semantic anchoring across textual and visual domains.
@@
[…]
Case Study #2: Multi-Agent Debate
Role-conditioned anchors demonstrate parallel semantic anchoring:
[…]
Each anchor concentrates the posterior over pattern classes on distinct argumentative regions. Both achieve $S(\mathcal{A}^{(\pm)})>\alpha_{c}$ due to high $\rho_{d}(P)$ for policy arguments and low $d_{r}$ from clear position specification. Since $k=0$ (instruction-only anchoring), both agents operate in the supercritical regime, ensuring reliable generation of position-specific arguments. The arbitration prompt activates synthesis patterns, demonstrating multi-anchor coordination within the unified framework. (Details in Appendix C.)
Case Study #3: Fine-Tuning Effects
Fine-tuning demonstrates pattern repository modification:
Medical Diagnosis Example:
  • Baseline: “Patient with fever, cough, fatigue” → “Could be cold, flu, or infections” (vague)
  • Fine-tuned: Same query → “Likely viral upper respiratory infection. Consider influenza A/B, COVID-19, rhinovirus. Recommend vital signs, rapid tests” (specific, actionable)
Fine-tuning increases $\rho^{\prime}_{d}(P_{\text{medical}})\gg\rho_{d}(P_{\text{medical}})$ and reduces $d_{r}^{\prime}(P_{\text{medical}},T)$, lowering the critical threshold $\alpha_{c}^{\text{fine-tuned}}<\alpha_{c}^{\text{original}}$ and enabling zero-shot success. This targeted improvement creates trade-offs: enhanced target domain performance but potential degradation in other areas, as demonstrated in our numeral base experiments.
Case Study #4: RAG Knowledge Integration
RAG demonstrates external pattern augmentation:
@@
  • RAG: Recent arXiv papers on quantum advantage, error correction.
  • Response: Synthesizes internal quantum fundamentals with recent breakthroughs.
RAG operates by augmenting pattern density: $\rho_{\text{total}}(P)=\rho_{d}(P)+\rho_{\text{ext}}(P)$, enabling threshold crossing for out-of-distribution queries where internal LLM knowledge provides insufficient anchoring strength.
@@
These case studies systematically confirm $\mathsf{UCCT}$’s core predictions: (1) threshold-crossing behavior manifests across textual, visual, and multi-agent domains, (2) the same anchoring strength formula governs diverse techniques, and (3) semantic anchoring represents a unified mechanism underlying few-shot learning, fine-tuning, RAG, and multi-agent reasoning.
Experiment: Validating $\mathsf{UCCT}$ Predictions
@@
Arithmetic tasks provide an ideal experimental testbed because they offer precise control over pattern familiarity through numeral base selection while maintaining constant computational complexity. Unlike domain-specific tasks where pattern density correlates with semantic difficulty, mathematical operations isolate representational factors from algorithmic complexity.
@@ 1. Design: Pattern Density Manipulation
A grep on CommonCrawl (_10 + $\gg$ _8 + $\gg$ _9 +) confirms this order-of-magnitude gap. Hence, we expect $\rho_{d}(10)>\rho_{d}(8)>\rho_{d}(9)$; The results reported in step 3 empirically verifies that ordering.
@@
For each base $B\in\{8,9,10\}$, we treat two-digit addition as a distinct latent pattern class $P_{B}$. Each sample includes explicit base tagging to prevent cross-base interference:
@@
Rank-16 adapters trained for one epoch on equation-answer pairs, primarily increasing base-specific pattern density $\rho_{d}(P_{B})$
@@
Same adapter training with four-step reasoning traces, reducing representational distance $d_{r}(P_{B},T)$ through procedural alignment
@@
Frozen backbone with $k$ prepended examples, directly testing the functional form $S(\mathcal{A})=\alpha\rho_{d}-\beta d_{r}-\gamma\log k$
2. Experimental Protocol
@@
For each base $B$ and shot count $k$ we prepend $k$ randomly drawn 2‑digit exemplars to the query, then measure accuracy on the $250$ ID‑2d problems. Plotting accuracy versus $k$ and fitting $\sigma(\alpha\rho_{d}-\beta d_{r}-\gamma\log k)$ yields (i) the 50% shot threshold $k_{50}$ and (ii) the phase‑width (10%–90% rise).
@@
Using the frozen encoder we take the final hidden state of the <eos> token as the embedding.
$\rho_{d}(B)$ $\;=\;\Bigl{[}\tfrac{1}{\binom{k}{2}}\sum_{i<j}\|a_{i}-a_{j}\|_{2}\Bigr{]}^{-1},$
$d_{r}(B)$ = $\;=\;1-\cos\bigl{(}T,\bar{A}\bigr{)},$
where $T$ is the query vector and $\bar{A}$ the centroid of $k\!=\!8$ anchor vectors (equation+answer only, no reasoning string).
These proxies capture the intuition that familiar patterns (high $\rho_{d}$) cluster tightly in representation space, while semantic alignment (low $d_{r}$) corresponds to high cosine similarity between query and context embeddings.
@@
3. Experimental Results
3.1 Pattern density hierarchy confirmed
@@
$(\rho_{d},\,d_{r})=\begin{cases}(12.69\pm 0.84,\;15.17\pm 1.23)&\text{base 10}\\ (9.67\pm 0.71,\;12.14\pm 0.98)&\text{base 8}\\ (9.62\pm 0.69,\;12.10\pm 1.01)&\text{base 9}\end{cases}$ (7)
The hierarchy $\rho_{10}>\rho_{8}\approx\rho_{9}$ confirms distinct pattern density regimes. Base-10’s substantially higher density (Cohen’s $d>2.8$ vs. other bases) reflects extensive decimal exposure in pre-training, while bases 8 and 9 show statistically indistinguishable densities despite different frequencies.
3.2 Few-shot phase transitions validate threshold predictions
@@
Table 1: Few-shot learning statistics from sigmoid fits. The monotonic ordering in all metrics ($10<8<9$) supports $\mathsf{UCCT}$’s prediction that learning efficiency scales as $\rho_{d}/d_{r}$.
The 10-fold difference between base-10 and base-9 thresholds ($k_{50}:0.28$ vs. $2.91$ shots) demonstrates the profound impact of pattern density on learning efficiency. Phase widths follow the same ordering, indicating that high-density patterns not only anchor faster but also exhibit sharper learning transitions.
3.3 Cross-base interference validates hysteresis predictions
Fine-tuning experiments reveal asymmetric interference patterns that validate $\mathsf{UCCT}$’s memory hysteresis predictions. We train base-specific LoRA adapters and measure cross-base accuracy changes:
@@ 3.3 Cross-base interference validates hysteresis predictions
Key findings:
  • Low-density base tuning: Training on base-8 or base-9 causes substantial accuracy drops in other bases ($\Delta=-15.3\%$ to $-28.7\%$).
  • High-density base tuning: Training on base-10 preserves performance on bases 8 and 9 ($\Delta=-2.1\%$ to $-4.6\%$).
  • Supervision format effects: Chain-of-thought reduces interference by $\sim$40through shared procedural structure.
This asymmetry validates $\mathsf{UCCT}$’s hysteresis prediction: high-density patterns remain supercritical under moderate perturbation, while low-density patterns are easily pushed subcritical by competing anchors.
3.4 Scope generalization validates semantic distance dynamics
@@
Figure 4: Scope generalization accuracy changes after fine-tuning. While 2-digit ID performance improves (positive bars), longer operands often show degradation, especially under CoT supervision that has rigid format expectations.
Chain-of-thought supervision exhibits a clear trade-off: while boosting 2-digit accuracy (+8.3% average), it reduces longer-operand performance (-12.7% for 4-digit problems). This validates $\mathsf{UCCT}$’s prediction that supervision methods can increase effective semantic distance $d_{r}$ for out-of-scope queries by creating overly specific procedural templates.
4. Experimental Summary: $\mathsf{UCCT}$ Validation
The arithmetic experiments provide comprehensive empirical validation of the $\mathsf{UCCT}$ framework in all three core predictions:
@@ 4. Experimental Summary: $\mathsf{UCCT}$ Validation
The relationship $k_{50}\propto d_{r}/\rho_{d}$ receives strong empirical support. Base-10’s 10× pattern density advantage translates directly to 10× fewer required examples, demonstrating that representational familiarity fundamentally determines learning efficiency. The monotonic ordering across all metrics ($k_{50}$, phase width, final accuracy) validates $\mathsf{UCCT}$’s quantitative predictions.
@@
The functional form $S(\mathcal{A})=\alpha\rho_{d}-\beta d_{r}-\gamma\log k$ accurately governs performance across different anchoring methods. Few-shot learning, LoRA fine-tuning, and chain-of-thought supervision all follow the predicted relationships between pattern density, semantic distance, and context length.
@@
Task-specific fine-tuning shifts pattern density and semantic distance terms in predictable ways, explaining ID gains, OOD drops, and cross-base interference. The asymmetric interference patterns validate $\mathsf{UCCT}$’s memory hysteresis predictions: high-density patterns resist perturbation while low-density patterns are vulnerable to competing anchors.
@@
Beyond qualitative insights, $\mathsf{UCCT}$ provides accurate quantitative predictions. The 10-fold threshold differences, asymmetric interference magnitudes ($\Delta=-15.3\%$ to $-28.7\%$ vs. $-2.1\%$ to $-4.6\%$), and scope generalization trade-offs all align with theoretical expectations, demonstrating the framework’s predictive power.

### v4

@@ 1.1.1 Redefining Statistical Priors
Without anchors, all LLMs default to the maximum likely prior pattern $P_{\text{prior}}$, subtraction.
Two-shot anchor redefining “-” as addition:
Example 1: 2 - 3 = 5  Example 2: 7 - 4 = 11  Question: 15 - 8 = ?  Answer: 23  (100% reinterpretation as addition)
Despite extensive prior training on subtraction, just two clear demonstrations override $P_{\text{prior}}$. This validates Equation 1, where $S(\mathcal{A})$ exceeds $S_{c}$ due to high $\rho_{d}(P_{T})$ and low $d_{r}(P_{\text{prior}},P_{T})$.
Ambiguous Anchors Near the Threshold:
Example 1: 33 - 27 = 60  Example 2: 11 - 9 = 20  Prompt: 15 - 8 = ?
  • Claude: $|a-b|\times 10$ (→ 70)
  • GPT-4o: $(a-b)\times 10$ (→ 70)
  • DeepSeek, Gemini: $a+b$ (→ 23)
These divergent outputs reflect marginal anchoring strength near $S_{c}$, where model-specific priors influence the posterior pattern selection. (See Appendix B for detailed breakdown.)
1.1.2 Visual Anchoring via Few-Shot Cat Classification
Figure 1 demonstrates that $\mathsf{UCCT}$’s threshold-crossing principles apply across modalities. The 4-shot anchoring aligns with the theory: $P_{\text{prior}}$ provides latent patterns, labeled cat images establish $\rho_{d}(P_{T})$, and $S(\mathcal{A})>S_{c}$ triggers a phase transition to reliable classification. This validates the theory’s cross-modal applicability.
@@ 1.1.2 Visual Anchoring via Few-Shot Cat Classification
Figure 1: UCCT Insight: Intelligence emerges from unconscious patterns + conscious anchoring. Top: few-shots (right) match patterns in the repository (left), yielding the association of cat $\rightarrow$ {p1, p2, p3, p5}. Bottom: test image activates its pattern {p1, p2, p6} and computes the overlap with the association, resulting pp(test image = cat) = high.
1.2 Role-Based Anchoring in Multi-Agent Debate
Role-conditioning shows parallel anchoring under instruction-only conditions ($k=0$):
$\mathcal{A}^{(+)}$ $=\text{``Support nuclear power; provide arguments.''}$
$\mathcal{A}^{(-)}$ $=\text{``Oppose nuclear power; provide arguments.''}$
Each anchor steers the posterior toward distinct argumentative regions. Both achieve $S_{(\pm)}>S_{c}$ via high $\rho_{d}(P_{T})$ and low $d_{r}(P_{\text{prior}},P_{T})$. Arbitration prompts activate synthesis patterns, illustrating multi-anchor coordination under the same probabilistic model. (See Appendix C for transcripts.)
1.3 Posterior Density Amplification via RAG
RAG increases posterior density w/o modifying the prior:
@@ 1.3 Posterior Density Amplification via RAG
  • RAG: Fetches recent arXiv papers on quantum advances.
  • Response: Merges internal priors with retrieved content.
RAG boosts $\rho_{d}(P_{T})$ via external documents: $\rho_{\text{total}}(P_{T})=\rho_{d}(P_{T})+\rho_{\text{ext}}(P_{T})$. This enables threshold-crossing for out-of-distribution queries without altering $P_{\text{prior}}$.
@@
These demonstrations confirm all three $\mathsf{UCCT}$ predictions: (1) threshold-crossing governs behavior, (2) a shared formula underlies all adaptation types, and (3) few-shot, RAG, and multi-agent settings operate via competition between $P_{\text{prior}}$ and $P_{T}$.
Part 2: Arithmetic Pattern Density and Threshold Validation
@@
Arithmetic tasks provide an ideal experimental testbed because they offer precise control over pattern familiarity through numeral base selection while maintaining constant computational complexity, as also seen in standard math benchmarks like GSM8K and MATH (Cobbe et al. 2021; Hendrycks et al. 2021a). Unlike domain-specific tasks where pattern density correlates with semantic difficulty, mathematical operations isolate representational factors from algorithmic complexity.
@@ 1. Design: Pattern Density Manipulation
A grep on CommonCrawl (_10 + $\gg$ _8 + $\gg$ _9 +) confirms this order-of-magnitude gap. Hence, we expect $\rho_{d}(10)>\rho_{d}(8)>\rho_{d}(9)$; the result reported in 2.3.1 empirically verifies that ordering.
@@
For each base $B\in\{8,9,10\}$, we treat two-digit addition as a distinct latent pattern class. The prior patterns $P_{\text{prior}}^{(B)}$ reflect pre-training exposure to base-$B$ arithmetic, while posterior patterns $P_{T}^{(B)}$ emerge from anchoring examples. Each sample includes explicit base tagging to prevent cross-base interference:
@@
Rank-16 adapters trained for one epoch on equation-answer pairs, reshaping the prior $p(P|C)$ and primarily increasing base-specific posterior pattern density $\rho_{d}(P_{T}^{(B)})$
@@
Same adapter training with four-step reasoning traces, reducing representational distance $d_{r}(P_{\text{prior}}^{(B)},P_{T}^{(B)})$ through procedural alignment
@@
Frozen backbone with $k$ prepended examples, directly testing the functional form $S(\mathcal{A})=\rho_{d}(P_{T})-d_{r}(P_{\text{prior}},P_{T})+\log k$
2.1 Experimental Protocol
@@
For each base $B$ and shot count $k$ we prepend $k$ randomly drawn 2‑digit exemplars to the query, then measure accuracy on the $250$ ID‑2d problems. Plotting accuracy versus $k$ and fitting $\sigma(\rho_{d}(P_{T})-d_{r}(P_{\text{prior}},P_{T})+\log k)$ yields (i) the 50% shot threshold $k_{50}$ and (ii) the phase‑width (10%–90% rise).
@@
Pattern density and semantic distance require operationalization through model representations. Using the frozen encoder, we extract final hidden states from the <eos> token position to compute prior-posterior competition metrics:
$\rho_{d}(P_{T})$ $\;=\;\Bigl{[}\tfrac{1}{\binom{k}{2}}\sum_{i<j}\|a_{i}-a_{j}\|_{2}\Bigr{]}^{-1},$
$d_{r}(P_{\text{prior}},P_{T})$ = $\;=1-\cos\bigl{(}\mathbf{e}_{\text{prior}},\mathbf{e}_{T}\bigr{)},$
where $\mathbf{e}_{\text{prior}}$ is the zero-shot query embedding (revealing prior patterns) and $\mathbf{e}_{T}$ is the centroid of k=8k\!=\!8 anchor example embeddings (revealing posterior patterns).
These proxies capture the $\mathsf{UCCT}$ intuition: familiar posterior patterns (high $\rho_{d}$) cluster tightly in representation space, while semantic alignment (low $d_{r}$) corresponds to high cosine similarity between prior and posterior anchor patterns.
@@
2.3 Experimental Results
2.3.1 Pattern density hierarchy confirmed
@@
$(\rho_{d}(P_{T}),\,d_{r}(P_{\text{prior}},P_{T}))=\begin{cases}(12.69\pm 0.84,\;15.17\pm 1.23)&\text{B10}\\ (9.67\pm 0.71,\;12.14\pm 0.98)&\text{B8}\\ (9.62\pm 0.69,\;12.10\pm 1.01)&\text{B9}\end{cases}$ (7)
The hierarchy $\rho_{10}>\rho_{8}\approx\rho_{9}$ confirms distinct posterior pattern density regimes. This supports the $\mathsf{UCCT}$ hypothesis that exposure-driven density directly lowers the anchor threshold required for adaptation. The similarity between bases 8 and 9, despite differing corpus frequencies, suggests threshold effects in posterior pattern formation.
2.3.2 Few-shot phase transitions validate threshold predictions
@@
Table 1: Few-shot learning statistics from sigmoid fits. The monotonic ordering in all metrics ($10<8<9$) supports $\mathsf{UCCT}$’s prediction that learning efficiency scales as $\rho_{d}(P_{T})/d_{r}(P_{\text{prior}},P_{T})$.
The 10-fold difference between the base-10 and base-9 thresholds ($k_{50}:0.28$ vs. $2.91$ shots) demonstrates the profound impact of the density of prior-posterior pattern on learning efficiency. Phase widths follow the same ordering, indicating that high-density posterior patterns not only anchor faster but also exhibit sharper learning transitions.
2.3.3 Cross-base interference validates hysteresis predictions
Fine-tuning experiments reveal asymmetric interference patterns that validate $\mathsf{UCCT}$’s memory hysteresis predictions. We train base-specific LoRA adapters and measure cross-base accuracy changes with three findings:
@@ 2.3.3 Cross-base interference validates hysteresis predictions
  • Low-density base tuning: Training on base-8 or base-9 modifies the prior $p(P|C)$ in ways that cause substantial accuracy drops in other bases ($\Delta=-15.3\%$ to $28.7\%$), as the new prior patterns interfere with cross-base posterior formation.
  • High-density base tuning: Training on base-10 preserves performance on bases 8 and 9 ($\Delta=-2.1\%$ to $4.6\%$), as the robust prior maintains sufficient density for cross-base anchoring.
  • Supervision format effects: Chain-of-thought reduces interference by $\sim$40% through shared procedural structure that decreases $d_{r}(P_{\text{prior}},P_{T})$ across bases.
This asymmetry validates $\mathsf{UCCT}$’s hysteresis prediction: high-density prior patterns remain supercritical under moderate perturbation, while low-density patterns are easily pushed subcritical by competing posterior anchors.
2.3.4 Scope generalization validates semantic distance dynamics
@@
Figure 4: Scope generalization accuracy changes after fine-tuning. While 2-digit ID performance improves (positive bars), longer operands often show degradation, especially under CoT supervision that increases $d_{r}(P_{\text{prior}},P_{T})$ for out-of-scope queries.
Chain-of-thought supervision exhibits a clear trade-off: while boosting 2-digit accuracy (+8.3% average), it reduces longer-operand performance (-12.7% for 4-digit problems). This validates $\mathsf{UCCT}$’s prediction that supervision methods can increase effective semantic distance $d_{r}(P_{\text{prior}},P_{T})$ for out-of-scope queries by creating overly specific procedural templates that increase the gap between prior patterns and posterior requirements.
2.4 Experimental Summary: $\mathsf{UCCT}$ Validation
The arithmetic experiments provide strong empirical validation of the $\mathsf{UCCT}$ framework on its three core predictions:
@@ 2.4 Experimental Summary: $\mathsf{UCCT}$ Validation
The relationship $k_{50}\propto d_{r}(P_{\text{prior}},P_{T})/\rho_{d}(P_{T})$ is strongly supported. Base-10’s 10× posterior pattern density advantage yields 10× fewer required examples, showing that prior-posterior representational familiarity governs learning efficiency. The monotonic ordering across metrics ($k_{50}$, phase width, final accuracy) confirms $\mathsf{UCCT}$’s quantitative predictions.
@@
The form $S(\mathcal{A})=\rho_{d}(P_{T})-d_{r}(P_{\text{prior}},P_{T})+\log k$ accurately captures performance across anchoring methods. Few-shot learning, LoRA fine-tuning, and chain-of-thought supervision align with predicted relationships between posterior density, prior-posterior distance, and context length.
@@
Task-specific fine-tuning alters prior pattern distributions $p(P|C)$ predictably, explaining ID gains, OOD drops, and cross-base interference. Asymmetric interference patterns validate $\mathsf{UCCT}$’s memory hysteresis: high-density priors resist perturbation; low-density priors are vulnerable.
@@
$\mathsf{UCCT}$ offers accurate quantitative predictions. The 10-fold threshold shifts, asymmetric interference ($\Delta=-15.3\%$ to $28.7\%$ vs. $2.1\%$ to $4.6\%$), and scope generalization trade-offs all match theoretical expectations.
Part 3: Geometric Trajectory Analysis
Objective.
To mechanistically ground $\mathsf{UCCT}$, we analyze how instruction and example representations evolve layer by layer in decoder-only LLMs (Meta-LLaMA-3.1-8B, Phi-4, Qwen3-14B) across 25 reasoning tasks. These tasks span various cognitive domains, including common sense reasoning (Talmor et al. 2019), logical inference (Liu et al. 2020), general knowledge reasoning (Clark et al. 2018), arithmetic and code synthesis (Hendrycks et al. 2021b).
Method.
For each transformer layer, we compute:
  • Pattern Density $\rho_{d}$: Inverse mean pairwise cosine distance among example embeddings, where each embedding is computed as the mean-pooled hidden states of tokens in the example span.
  • Mismatch $d_{r}$: Cosine distance between the centroids of the instruction and the example embeddings.
Preliminary Findings (More results in App D).
  • U-shaped $\rho_{d}$ trajectories support $\mathsf{UCCT}$’s prediction of three computational stages:
    1. 1.
      Enrichment: Early specialization reduces density.
    2. 2.
      Abstraction: Mid-layer reasoning aligns examples with instructions ($d_{r}$ dips).
    3. 3.
      Standardization: Final re-clustering prepares output structure.
  • Task difficulty tracks $d_{r}$. Abstract tasks (e.g., logic) exhibit higher $d_{r}$, indicating greater representational transformation is needed to align instructions and examples (Liu et al. 2020).
  • Cognitive styles differ. Despite shared U-shapeds, each model exhibits a distinct trajectory, reflecting stylistic divergence from architectural and training differences, consistent with $\mathsf{UCCT}$’s interpretation of model-specific priors.
Figure 5: Layer-wise evolution of pattern density $\rho_{d}$ and mismatch $d_{r}$ in LLaMA-3.1-8B. The U-shape reflects enrichment, abstraction, and standardization stages predicted by $\mathsf{UCCT}$. Full results across models are in App D.